Deleted Question

Deleted Question
Answer #1

total acceleration with a magnitude of 15.0 m/s2 and a constant tangential acceleration of 12.0 m/s2

So radial acceleration=sqrt(15^2-12^2)=9 m/s^2=v^2/r

v=sqrt(36.9)=6.07 m/s

after 1/8 seconds,

Only tangentional accelartion changes the speed,radial will change only direction

v=u+at

=6.07+12t

=7.57 m/s